OpenStudy (anonymous):

16^(6x) * (1/4)^(9x-5) = 64 * (4^x)^-7 Solve

OpenStudy (anonymous):

are you kidding?

OpenStudy (anonymous):

Why would I be?

OpenStudy (anonymous):

gimmick is to rearrange this so that both sides are powers of 4. you get \[4^{3x-7}\] on the left

OpenStudy (anonymous):

sorry. that is the right hand side, not the left hand side

OpenStudy (anonymous):

right hand side is \[4^{3x-7}\]

OpenStudy (anonymous):

left hand side is \[4^{3x+5}\]

OpenStudy (anonymous):

once that algebra nightmare is over it is pretty simple. since they are both powers of 4 you can just set the exponents equal and solve: \[3x+5=3-7x\] \[10x=-2\] \[x=\frac{-2}{10}=-\frac{1}{5}\]

OpenStudy (anonymous):

oh lord i see i made a typo in earlier post sorry

OpenStudy (anonymous):

right hand side is \[4^{3-7x}\] not what i wrote

OpenStudy (anonymous):

answer is correct though

OpenStudy (anonymous):

No problem, thanks! I see where I went wrong XD

OpenStudy (anonymous):

hope you see where i went wrong too. but answer is correct.

OpenStudy (anonymous):

\[\color{blue}{\text{hello my myininaya!}}\]

myininaya (myininaya):

im sorry but are you sure you wrote the problem down correctly autumn

OpenStudy (anonymous):

\[\color{#FF0018}{\text{ are you in the pink?}}\]

OpenStudy (anonymous):

\[\color{red}{\text{don't cause trouble}}\]

myininaya (myininaya):

hmmm maybe i'm wrong i read the last thingy as x^(-7)

OpenStudy (anonymous):

\[16^{6x} \times( \frac{1}{4})^{9x-5} = 64 \times (4^x)^{-7}\]

myininaya (myininaya):

okay that looks prettier :)

myininaya (myininaya):

i accept the answer as -.2 now

OpenStudy (anonymous):

as i said my little \[\color{red}{\text{apple pi}}\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

right hand side is \[4^{7-3x}\]

OpenStudy (anonymous):

ok maybe it is is \[4^{3-7x}\]

OpenStudy (anonymous):

\[\color{blue}{\text{typo!}}\]

myininaya (myininaya):

it is the second one

OpenStudy (anonymous):

what kind of chump math teacher would give this problem?

OpenStudy (anonymous):

stupid gimmick of writing everything in the same base. sheesh